From the following table calculate for the year 2019:

  1. Capital
  2. Retained Earnings
  3. Equity

2018

2019

Beginning Capital

$100,000

Issued Capital

50,000

$20,000

Treasury Stock

10,000

6,000

Beginning Retained Earnings

25,000

Net Income

10,000

7,000

Dividend Paid

5,000

4,000

Can someone help me with this question, for the Financial Statement Analysis course?
Answer #1
2018
2018
YEAR 2018
A Beginning Capital $100,000
B Issued Capital 50,000
C Treasury Stock 10,000
D=A+B-C Ending Capital 140,000
E Beginning Retained Earnings 25,000
F Net Income 10,000
G Dividend Paid 5,000
H=E+F-G Ending Retained Earnings 30,000
2019
YEAR 2019 2019
A Beginning Capital 140,000
B Issued Capital $20,000
C Treasury Stock 6,000
D=A+B-C Ending Capital $154,000
E Beginning Retained Earnings 30,000
F Net Income 7,000
G Dividend Paid 4,000
H=E+F-G Ending Retained Earnings 33,000
Ending Equity =154000+33000 $187,000
2018 2019
Beginning Capital $100,000 140,000
Issued Capital 50,000 $20,000
Treasury Stock 10,000 6,000
Beginning Retained Earnings 25,000 30,000
Net Income 10,000 7,000
Dividend Paid 5,000 4,000